one of the angles of a parallelogram is 40 degrees. what is the measure of one of the obtuse angles? why?
: one of the angles of a parallelogram is 40 degrees. what is the measure of one of the obtuse angles? why?

Answer by jojo14344(1031)

Well, the opposite angle should be another 40 degrees. Now the obtuse angle should be 140 degrees, of course likewise opposite of it is 140 degrees too.
One property of a parallelogram being that opposite angles are congruent.
Why 140 degrees?
Let's say ABCD is a parallelogram. Angle A is 40 degrees and opposite which is angle C is also 40 deg. Now outside of angle A is 140deg (bec. 140+40=180) and that 140 deg is equal to angle B which is congruent to angle D.Therefore, angle B=140deg and angle D= 140deg
The sum of all included angles should be 360 degrees, which is right.
A+B+C+D=40+140+40+140=360deg
Thank you,
Jojo